An adiabatic air compressor compresses 100 L/S of air at 120 kPa and 20°C to 1000 kPa and 300°C. Determine (a) the work required by the compressor, in kJ/kg, and (b) the power required to drive the air compressor, in kW.
Answer #1

given P=120 kPa P=200 = 293 * a diabetic - Po = 2000 kPa - 30 0 C - 513k V- 100 Ysec = 10 m/s * P, = PORT 120 = P2O. 2874293please like